The grade point averages for 10 randomly selected high school students are listed below. Assume the grade point averages are normally distributed. ​2.0, 3.2,​ 1.8, 2.9,​ 0.9, 4.0,​ 3.3, 2.9,​ 3.6, 0.8 Find a​ 98% confidence interval for the true mean.
Answer #1

Solution :

We have,

= 25.7 / 10 = 2.57

s = 1.145

n = 10

df = 10 - 1 = 9

t /2,df = 2.896

Margin of error = E = t/2,df * (s /n)

= 2.896 * (1.145 / 10)

Margin of error = E = 1.05

The 98% confidence interval estimate of the population mean is,

- E < < + E

2.57 - 1.05 < < 2.57 + 1.05

1.52 < < 3.62

(1.52 , 3.62)